2. (2) PTS) The euro and the dollar uro Suppose that on January Ist 2017 the euro exchange rate was 1,05 US dollars per e 05), the one year interest rate in the USA was 2% and that in the Euro zone was 3.5%. If the uncovered parity condition holds true, a. What are markets expecting for the exchange rate to happen over the year 2017? When will the interest parity condition not hold true? b. If the exchange rate on January It 2018 ends up being equal to 1,1, what would have been the best choice for an investment fund to make on January 2017 for the rest of the year 2017? c. Starting from the initial point of part a) above, suppose a new US expansionary fiscal policy (Trump) not accommodated by the US monetary authority, what would be the impact on the current exchange rate dollar-euro? Could it affect also the expected exchange rate for next year? In what direction? Suppose the foreign interest rate rises by one percentage point to 3%. d. What are the main determinants of the exchange rate in a small open economy? What about the determinants of the real exchange rate? Explain your results, use graphies when appropriate
Answer #1

Uncovered interest rate parity is the condition in which the difference in interestr ates between two countries is equal to the expected change in exchange rates between those countries' currencies.

a)

Interest rate in USA iU =2%

Interest rate in Euro Zone iE = 3.5%

Therefore % change in Exchange rate should be 2% -3.5% = -1.5%

So exchange rate from USD to Eur should decrease by 1.5%

or new exchange rate should be 1.05*(1-1.5%) = 1.05 * 98.5% = 1.0345%

As per uncovered interest parity, Euro to USD exchange rate should get lower and converge at a point where there will be no arbitrage opportunity. That is change in exchange rate and difference in interst rate of two countries should be same so that investor don't generate any return. The interest parity condition will not hold true when USD will depreciate in terms of euro. As per parity condition interest rate and exchange rate should move in same direction not oppossite.

b) If 1 euro became equal to 1.1 US dollar from 1.05 US dollar, investor should have followed following strategy. As euro is appreciating in terms of US, investor should have borrowed US dollars and would have bought Euro dollars from it. Then after one year the investor's capital in USD would have increased (1.10/1.05) times or .1.0476 times.

c) Expansionary fiscal policies may result excessive fiscal deficit, which may lead to printing of more currency notes which will in turn depreciates local currency. Also due to expansionary fiscal policy, disposable income of people will increase which will in turn increase imports and hence increases the demand of currency which in turn may lead to depreciation of local currency.

So based on these assumptions, US dollar is expected to depreciate in comparison to Euro. Also since interest rate is increasing from 2% to 3%, difference between two interest rate will remain only .5%.

So next year exchnage rate will be 1.05 * 99.5% = 1.04475 which would have been 1.0345% at 2% interest rate.

d) Determinants of exchange rate in small open economy are as follows:

  1. Difference in inflation rate of two countries
  2. Difference in interest rate of two countries.
  3. Current Account Deficits

  4. PublicDebt

  5. Terms of Trade (change in export and import)

  6. Political Stability and Economic Performance and of both countries

Determinants of real exchange rates are money supply, domestic and foreign interest rate, real GDP, real government expenditure, deficit per GDP, domestic and foreign outstanding debt per GDP, domestic and foreign externally financed debt per GDP and commodity price.